Find the solution of


(D^4+2D^3-6D^2-16D-8)y=0

1
Expert's answer
2022-02-08T12:09:15-0500

Solution:

"(D^4+2D^3-6D^2-16D-8)y=0"

"\Rightarrow D^4+2D^3-6D^2-16D-8=0
\\\Rightarrow \left(D+2\right)^2\left(D^2-2D-2\right)=0
\\ \Rightarrow D+2=0,D+2=0\quad \mathrm{or}\quad \:D^2-2D-2=0
\\\Rightarrow D=-2,D=-2,\:D=1+\sqrt{3},\:D=1-\sqrt{3}"

Thus, solution will be:

"y=c_1e^{\left(1+\sqrt{3}\right)x}+c_2e^{\left(1-\sqrt{3}\right)x}+c_3e^{-2x}+c_4xe^{-2x}"